### v1

Abstract
Weather forecasting is a crucial task for meteorologic research, with direct social and economic impacts. Recently, data-driven weather forecasting models based on deep learning have shown great potential, achieving superior performance compared with traditional numerical weather prediction methods. However, these models often require massive training data and computational resources. In this paper, we propose EWMoE, an effective model for accurate global weather forecasting, which requires significantly less training data and computational resources. Our model incorporates three key components to enhance prediction accuracy: meteorology-specific embedding, a core Mixture-of-Experts (MoE) layer, and two specific loss functions. We conduct our evaluation on the ERA5 dataset using only two years of training data. Extensive experiments demonstrate that EWMoE outperforms current models such as FourCastNet and ClimaX at all forecast time, achieving competitive performance compared with the state-of-the-art Pangu-Weather model in evaluation metrics such as Anomaly Correlation Coefficient (ACC) and Root Mean Square Error (RMSE). Additionally, ablation studies indicate that applying the MoE architecture to weather forecasting offers significant advantages in improving accuracy and resource efficiency.
keywords:
Weather Forecasting, Deep Learning, Mixture-of-Experts, ERA5
1 Introduction
Weather forecasting is the analysis of past and present weather observations, as well as the use of modern science and technology, to predict the state of the Earth atmosphere in the future. It is one of the most important applications of scientific computing and plays a crucial role in key sectors such as transportation, logistics, agriculture, and energy production [1]. Traditionally, atmospheric scientists have relied on Numerical Weather Prediction (NWP) methods [2, 3], which utilize mathematical models of the atmosphere and oceans to forecast the weather states based on current weather conditions. While modern meteorological forecasting systems have achieved satisfactory results using NWP methods, these methods largely rely on parametric numerical models, which can introduce errors in the parameterization [4] of complex, unresolved processes. Additionally, NWP methods face challenges in meeting the diverse needs of weather forecasting due to its high computational cost, the difficulty of solving nonlinear physical processes, and model deviations [5, 6].
To address the above issues of NWP models, researchers have turned their attention to data-driven weather forecasting based on deep learning methods. These methods run very quickly and can easily achieve a balance among model complexity, prediction resolution, and prediction accuracy [7, 8, 9]. Denby [10] first employed Convolutional Neural Network (CNN) for the classification of weather satellite images. Xu et al. [11] utilized a combination of Generative Adversarial Network (GAN) and Long Short-Term Memory (LSTM) for cloud prediction. While these attempts reveal the potential of deep learning methods in weather forecasting, they are limited by low-resolution data and ineffective models, resulting in limited applications. Recently, FourCastNet [12] increased the resolution to 0.25, comparable to the ECMWF Integrated Forecast Systems (IFS). ClimaX [13] showed superior performance on weather benchmarks for weather forecasting and climate projections, even when pretrained at lower resolutions and with limited computing budgets. Pangu-Weather [14] was the first state-of-the-art model to outperform IFS. These models are based on Vision Transformer (ViT) [15], and use standard ViT embedding to process meteorological data. However, meteorological data is different from general computer vision image input. The channels in meteorological data represent atmospheric variables with intricate physical relationships and have different coordinate information in the earth coordinate system. ViT embedding method cannot effectively extract the physical features between these meteorological variables and the geographical features of the variables themselves [16]. Moreover, although these innovative approaches yielded significant improvement, their adoption was constrained by the high computational demands required [17] for training. For example, FourCastNet utilized 64 Nvidia A100 GPUs for a training period of 16 hours, highlighting the extensive resources needed [18, 19] for the development of cutting-edge, deep learning based weather forecasting model.
In this work, we present an effective data-driven model called EWMoE for global weather forecasting. We start with a Vision Transformer (ViT) architecture and, to address the issues of deep learning based models, our EWMoE consists of three novel components: (1) meteorology-specific embedding, including patch embedding and position embedding, that fully models the interactions among atmospheric variables and their geographical location features; (2) a crucial Mixture-of-Experts (MoE) layer that increases the model capacity without increasing compute requirements, greatly improving model prediction accuracy with significantly less computational resources and training data; (3) an auxiliary loss and a position-weighted loss that perform specific optimizations during the model training process. We trained our proposed EWMoE on two years of data from the ERA5 dataset [20] using only 2 Nvidia 3090 GPUs in less than 9 days. Experiment results show that EWMoE significantly outperforms FourCastNet and Climax, and achieves a comparable level of forecasting accuracy as Pangu-Weather for short-range forecasting (1-3 days). As the forecast time increases, EWMoE exhibits more stable and excellent prediction results compared with them. Notably, EWMoE achieves this performance by training on less data and requiring orders-of-magnitude fewer GPU hours. Finally, we conduct extensive ablation studies to analyze the importance of individual components in EWMoE, demonstrating its potential for facilitating future works.
Overall, our contributions can be summarized as follows:
  • We propose EWMoE, an effective weather model with MoE for global weather forecasting, which demonstrates superior performance over other state-of-the-art models for short-to-medium-range weather prediction.
  • Our EWMoE consists of three main components: (1) a meteorology-specific embedding layer to fully extract the inherent information of weather data; (2) an MoE layer to increase the model capacity without increasing compute requirements; (3) two loss functions to optimize the training process.
  • Unlike other deep learning based models, EWMoE achieves these superior global weather predictions with significantly smaller number of computational resources and training data.
2 Related work
2.1 Numerical weather prediction
Numerical Weather Prediction (NWP) is a method used to forecast atmospheric conditions and weather states by utilizing systems of partial differential equations [1, 21, 22]. These equations describe different physical processes and thermodynamics, which can be integrated over time to obtain future prediction results. Although NWP models have good reliability and accuracy in weather forecasting, they face many challenges such as systemic errors [4, 23] produced by parametrization schemes. NWP methods also involve high computation costs due to the complexity of integrating a large system of partial differential equations [24], especially when modeling at high spatial and temporal resolutions. Furthermore, more observation data does not improve NWP forecast accuracy since models rely on the expertise of scientists in the meteorological field to refine equations, parameterizations and algorithms [25]. In recent years, many efforts have been made to improve the accuracy and efficiency of NWP models. For example, some researchers [26] have proposed grid refinement techniques to increase the model resolution, while others [27] have suggested fine-tuning physical parameterizations to further enhance the accuracy of weather forecasts.
2.2 Deep learning based weather forecasting
In order to address the challenges of NWP models, researchers have shown increasing interest in the application of deep learning models to weather forecasting [28, 29]. These models train deep neural networks to predict future weather states using vast amounts of historical meteorological data [30, 31, 32], such as the ERA5 reanalysis dataset. Compared with the traditional NWP models, deep learning based models have the potential to generate more accurate weather forecasts with less computational cost [33, 18, 34, 35]. Once trained, these models can produce timely forecast in a few seconds, which is considerably faster than NWP models that take hours or even days [36]. Weyn et al. [9] proposed an elementary weather prediction model using deep Convolutional Neural Networks (CNNs) trained on past weather data, although their method only achieves a modest resolution of 2.5 and contains no more than three variables per grid. However, rapid progress has been made in recent years. FourCastNet [12], a data-driven weather forecasting model, utilized the Vision Transformer (ViT) architecture and Adaptive Fourier Neural Operators (AFNO) [37], first pushing model resolution to 0.25 as NWP methods can. FourCastNet’s predictions are comparable to the IFS model at lead times of up to three days, pointing to the enormous potential of data-driven modeling in complementing and eventually replacing NWP. ClimaX [13] is the first model that can effectively scale using heterogeneous climate datasets during pretraining and generalize to diverse downstream tasks during fine-tuning, paving the way for a new generation of deep learning models for Earth systems science. Pangu-Weather [14], a 3D Earth-specific Transformer model, is the first to outperform operational Integrated Forecasting System (IFS) [38], producing even more favorable evaluation results on many of nine variables and pressure levels presented.
3 Preliminaries
3.1 Dataset
ERA5 [20] is a publicly available atmospheric reanalysis dataset produced by the European Centre for Medium-Range Weather Forecasts (ECMWF). It provides comprehensive information about the Earth climate and weather conditions, and is widely used in climate research, climate change analysis, weather forecasting, environmental monitoring, and other fields. The dataset covers the period from 1940 to the present and includes a wide range of meteorological variables such as temperature, humidity, wind speed, precipitation, cloud cover, and more. The spatial resolution of the data is 0.25 latitude and longitude, with hourly intervals, and 37 vertical pressure levels ranging from 1000 hPa to 1 hPa.
In this study, we use the ERA5 reanalysis dataset as the ground-truth for the model training, which has a spatial resolution of 0.25 (721$\times$1440 latitude-longitude grid points). Specifically, we select six-hourly sampled data points (T0, T6, T12, T18), with each sample consisting of twenty atmospheric variables across five vertical levels (see Table 1 for more details). In addition, to demonstrate the effectiveness of our model in the case of limited data and computing resources, we use two years of data for training (2015 and 2016), one year for validation (2017), and one year for testing (2018).
Table 1: The abbreviations are as follows: $U_{10}$ and $V_{10}$ represent the zonal and meridonal wind velocity from the surface, specifically, at a height of 10m; $T_{2m}$ represents the temperature at 2m from the surface; $T$, $U$, $V$, $Z$ and $RH$ represent the temperature, zonal velocity, meridonal velocity, geopotential and relative humidity at specified vertical level; $TCWV$ represents the total column water vapor.
Vertical level Variables
Surface $U_{10},V_{10},T_{2m},sp,mslp$
10000 hPa $U,V,Z$
850 hPa $T,U,V,Z,RH$
500 hPa $T,U,V,Z,RH$
50 hPa $Z$
Integrated $TCWV$
3.2 Weather forecasting task
Given a dataset of historical weather data, the task of global weather forecasting is to forecast the future global atmosphere states based on the current atmosphere conditions [39]. Specifically, we denote the initial weather state as $X_{i}\in\mathbb{R}^{C\times H\times W}$, where $C$ represents the number of atmosphere variables or channels, $H$ and $W$ are the height and width, respectively. Our model aims to generate 8-day forecasts $\{\hat{X}_{i+1},\hat{X}_{i+2},\cdots,\hat{X}_{i+32}\}$ with a time interval of six hours. However, it is challenging to train the model to directly forecast the future weather state $\hat{X}_{T}=f_{\theta}(X_{i})$ for each target lead time $T$. Since the weather system is chaotic, forecasting the future weather directly for large $T$ is difficult [40, 41, 42]. Moreover, it requires training one network for each lead time, which can be computationally expensive when the dataset is very large. To avoid this issue, we train our model to produce forecasts in an autoregressive manner. For longer forecasts, we unfold the model by iteratively feeding its predictions back to the model as input, e.g., $\hat{X}_{i+1}=f_{\theta}(X_{i}),\hat{X}_{i+2}=f_{\theta}(X_{i+1}),\cdots,\hat{X}_{i+32}=f_{\theta}(X_{i+31})$.
4 Methods
In this section, we introduce EWMoE, an effective weather model with Mixture-of-Experts (MoE) for global weather forecasting, as illustrated in Figure 1. Our EWMoE consists of three main components, which include: 1) meteorology-specific embedding including patch embedding and position embedding; 2) the structure of the MoE layer; 3) the auxiliary loss and the position-weighted loss for model training optimization.
Figure 1: Overall architecture of the proposed EWMoE model. Based on the standard encode-decoder design [43], EWMoE first uses a meteorology-specific embedding layer to extract the feature embeddings of input weather images. Then, an MoE layer routes the tokens to top-$k$ experts and integrates the outputs by the gating network. Finally, the feature representation is used to reshape the model output.
4.1 Meteorology-specific embedding
The information contained in weather data is very different from the natural images used in computer vision tasks. The channels in weather data represent different meteorological variables, and there are complex physical relationships among these variables. For example, there is a close relationship between temperature and relative humidity, while temperature and pressure obey the ideal gas law and are positively correlated. Therefore, effectively extracting the internal relations between these meteorology variables is the key to accurate weather forecasting.
Meteorology-specific patch embedding. The standard patch embedding in ViT, which simply uses a linear layer to project a patch into a vector, may not be able to sufficiently extract the internal relationships between different meteorological variables. Therefore, we adopt a meteorology-specific patch embedding module for our architecture. We denote the input image as a high dimension tensor $X\in\mathbb{R}^{C\times H\times W}$, and the module divides the input image into a sequence of patches, where the size is $p\times p$. Each input patch of size $p^{2}$ is linearly embedded to a vector of dimension $D$, where $D$ is the chosen embedding size. This results in $C\times(H/p)\times(W/p)$ patches in total. Then, a learnable query vector is used to perform cross-attention operation at each spatial position to conduct the interactions between the meteorological variables of each channel. The cross-attention layer outputs a sequence of shape $(H/p)\times(W/p)$, significantly reducing the sequence length and lowering the computational cost. Moreover, the cross-attention layer allows the model to learn the internal relationship between meteorological variables of different channels, which strengthens the ability of our model to extract complex physical features.
Meteorology-specific position embedding. For global weather forecasting, each input token corresponds to an absolute position on the Earth’s coordinate system. More importantly, some meteorology variables are closely related to their absolute position. For example, geopotential height is closely related to the latitude, while the wind speed and temperature are closely related to height. In this situation, using relative position embedding or 1D/2D position embedding does not capture this intrinsic feature well. Therefore, we use a 3D absolute position embedding for meteorology-specific position embedding, taking the 3D position information (height, latitude and longitude) of the patch into account. Specifically, for each input $D$ dimensional vector, we train three sets of learnable position embedding vectors with dimension of $D/3$. Each set corresponds to the absolute position of a patch on the Earth’s coordinate system, which are altitude, longitude and latitude respectively. After concatenating these three sets of vectors, we obtain the final 3D absolute position embedding vector with a dimension of $D$.
4.2 The structure of the MoE layer
Following the meteorology-specific embedding, we leverage sparsely activated Mixture-of-Experts (MoE) in our EWMoE model, which allows increasing the model capacity (total number of available parameters) without increasing computing requirements (number of active parameters). There are $N$ encoder blocks in EWMoE and we replace the dense Feed-Forward Network (FFN) layer present in encoder with a sparse MoE layer, as shown in Figure 1. Each MoE layer consists of a collection of independent feed forward networks as the “experts”. A gating network then uses a softmax function to route the input tokens to the best-determined top-$k$ experts. This means that for each given input token, only a small number of experts are activated, giving our model more flexibility and capacity to complete complex weather forecasting tasks with strong performance.
Given $N$ experts and input token $x$, the output $y$ of the MoE layer can be written as follows:
${y=\sum_{i=1}^{N}{g_{i}(x)E_{i}(x)}},$ (1)
where $g_{i}(x)$ is the output of the $i$-th element of gating network and $E_{i}(x)$ is the output of the $i$-th expert network. According to the formulation above, we can save computation based on the sparsity of the output $g(x)$. When $g(x)$ is a sparse vector, only a few experts would be activated and updated by back-propagation during training. Wherever $g_{i}(x)=0$, we need not compute the corresponding $E_{i}(x)$.
Top-$k$ routing. We use top-$k$ routing to select the top ranked experts, keeping only the top-$k$ gate values while setting the rest to $\infty$- ∞ before taking the softmax function. Then, the following $g(x)$ can be formulated as:
${g(x)=Softmax(top-k(x\cdot W+\epsilon,k))},$ (2)
$Top-k(m,k)_{i}=\left\{\begin{aligned} &m_{i}&&{\rm if}\>m_{i}\>{\rm is\>in\>the\>top-}\>k\>{\rm elements\>of\>}m,\\ &-\infty&&{\rm otherwise},\end{aligned}\right.$ (3)
where $W$ is a trainable weight matrix and $\epsilon\sim\mathcal{N}(0,\frac{1}{e^{2}})$ is a Gaussian noise for exploration of expert routing ($e$ is the mathematical constant). When $k\ll N$, most elements of $g(x)$ would be zero so that our model can achieve greater capacity while using less computation. In the MoE layer, we train our model with $k=2,N=20$.
4.3 Loss function for model training optimization
Auxiliary loss for load balancing. In the MoE layer, we dispatch each token to $k$ experts. There is a phenomenon that most tokens may be dispatched to a small portion of experts, as the favored experts are trained more rapidly and thus are selected even more by the gating network. Such an unbalanced distribution would decrease the throughput of our model, and as most experts would not be fully trained, the flexibility and performance of the model would be reduced. To resolve this issue, we use a differentiable load balancing auxiliary loss instead of separate load-balancing and importance-weighting losses proposed by Shazeer et al. [44] for a balanced loading in routing. Given $E$ experts and a batch $B$ with $L$ tokens, the following auxiliary loss is added to the total model loss during training:
${l_{aux}=E\cdot\sum_{i=1}^{E}{h_{i}\cdot P_{i}}},$ (4)
where $h_{i}$ is the fraction of tokens dispatched to expert $i$:
${h_{i}=\frac{1}{L}\sum_{x\in B}\mathbbm{1}\{{\rm argmax}\>g(x)=i\}},$ (5)
and $P_{i}$ is the fraction of the router probability distributed for expert $i$:
${P_{i}=\frac{1}{L}\sum_{x\in B}g_{i}(x)}.$ (6)
The goal of the auxiliary loss is to achieve a balanced distribution. When we minimize $l_{aux}$, we can see both $h_{i}$ and $P_{i}$ would close to a uniform routing.
Position-weighted loss. In weather forecasting tasks, the correct prediction of near-surface atmospheric variables is crucial, which has the most societal impact on human activities. We use a weighting function based on the laws of physics to weigh variables near the surface higher and employ the latitude-weighted mean squared error as our objective function. Given the prediction $\hat{X}_{i+\Delta t}$ and the ground-truth $X_{i+\Delta}$, the loss is written as:
${\mathcal{L}=\frac{1}{CHW}\sum_{c=1}^{C}\sum_{i=1}^{H}\sum_{j=1}^{W}f(v)L(i)\left(\hat{X}_{i+\Delta t}^{cij}-X_{i+\Delta t}^{cij}\right)^{2}},$ (7)
in which $f(v)$ is the weight value of variable $v$, and $L(i)$ is the latitude-weighting factor:
${L(i)=\frac{{\rm cos}({\rm lat}(i))}{\frac{1}{H}{\textstyle\sum_{i^{{}^{\prime}}=1}^{H}{\rm cos}({\rm lat}(i^{{}^{\prime}}))}}},$ (8)
where lat$(i)$ is the latitude of the $i$-th row of the grid. The latitude weighting term accounts for the non-uniformity when we grid the round globe. The grid cells toward the equator have larger areas than the cells near the pole, and thus should be assigned more weights.
5 Results
We compare our EWMoE with other state-of-the-art weather forecasting models, and show the results on predicting multiple meteorological variables. We also provide visualization examples to demonstrate the superiority of EWMoE in global weather forecasting. Additionally, we conduct extensive ablation studies to analyze the the importance of each component in our model. Our implementation code is available at https://github.com/Tomoyi/EWMoE.
5.1 Evaluation metrics
Following the previous deep learning based methods, the accuracy of deterministic forecast is computed by two quantitative metrics, namely, the latitude-weighted Root Mean Square Error (RMSE) and latitude-weighted Anomaly Correlation Coefficient (ACC).
The latitude weighted ACC for a forecast variable $v$ at forecast time-step $l$ is defined as follows:
${{\rm ACC}(v,l)=\frac{{\textstyle\sum_{m,n}}L(m)\hat{X}_{pred}(l)[v,m,n]\hat{X}_{true}(l)[v,m,n]}{\sqrt{{\textstyle\sum_{m,n}}L(m)({\hat{X}_{pred}(l)[v,m,n]})^{2}{\textstyle\sum_{m,n}L(m)({\hat{X}_{true}(l)[v,m,n]})^{2}}}}},$ (9)
where ${\hat{X}_{pred/true}(l)[v,m,n]}$ represents the long-term-mean-subtracted value of predicted or true variable $v$ at the location denoted by the grid co-ordinates $(m,n)$ at the forecast time-step $l$. The long-term mean of a variable is just the mean value of it over a large number of historical samples in the training dataset. The long-term mean-subtracted variables ${\hat{X}_{pred/true}}$ represent the anomalies of those variables that are not captured by the long term values. $L(m)$ is the latitude weighting factor at the co-ordinate $m$ which is defined in Eq. (8). The latitude-weighted RMSE for a forecast variable $v$ at forecast time-step $l$ is defined by the following equation:
${{\rm RMSE}(v,l)=\sqrt{\frac{1}{NM}\sum_{m=1}^{M}\sum_{n=1}^{N}L(m){(X_{pred}(l)[v,m,n]-X_{true}(l)[v,m,n])}^{2}}},$ (10)
where $X_{pred/true}(l)[v,m,n]$ represents the value of predicted or true variable $v$ at the location denoted by the grid co-ordinates $(m,n)$ at the forecast time-step $l$.
Figure 2: Latitude-weighted ACC results of EWMoE and the baselines predicting four key variables $Z500$, $T2m$, $T850$ and $U10$ in 2018 (higher ACC is better).
Figure 3: Latitude-weighted RMSE results of EWMoE and the baselines predicting four key variables $Z500$, $T2m$, $T850$ and $U10$ in 2018 (lower RMSE is better).
5.2 Comparison with state-of-the-art models
We compare the forecast performance of EWMoE with FourCastNet, ClimaX and Pangu-Weather, three leading deep learning methods for global weather forecasting. Figures 2 and Figure 3 evaluate different methods on forecasting four key weather variables at lead time from 1 to 8 days in terms of ACC and RMSE, respectively. The results show that EWMoE has both higher ACC and lower RMSE than FourCastNet [12] and ClimaX [13] for all the variables analyzed. For short-range forecasting (1-3 days), EWMoE demonstrates a comparable level of forecasting accuracy as Pangu-Weather [14]. In addition, as the forecast time increases, significant improvement with our EWMoE is observed and EWMoE outperforms Pangu-Weather from day 3, demonstrating EWMoE’s remarkable ability and stability for short-to-medium-range weather forecasting.
Moreover, we also note that EWMoE achieves this strong performance with much less training data and computing resources compared with the baselines. We train our EWMoE on 2 years of training data, which is approximately 18$\times$ less data than FourCastNet and ClimaX’s 37 years of training data, and 120$\times$ less than that used for Pangu-Weather, which uses 39 years of training data with 13 pressure levels. The training of EWMoE was completed under 9 days on 2 Nvidia 3090 GPUs. In contrast, FourCastNet took 16 hours to train on 64 A100 GPUs, ClimaX took 7 days on 80 V100 GPUs, and Pangu-Weather required 64 days on 192 V100 GPUs. Our novel weather forecasting MoE model generates accurate forecasts with much less training data and computational cost, which will facilitate future works that build upon our proposed framework.
5.3 Visualization
We visualize the predicted results of EWMoE at lead days 1, 3, 7 for two variables, Z500 (geopotential at the pressure level of 500 hPa) and U10 (the 10m zonal wind velocity), and compare the results with the ERA5 ground-truth. The initial time point is 00:00 UTC, January 15th, 2018. In Figure 4 and Figure 5, the first column shows the ERA5 ground-truth at that lead day, the second column shows the prediction result, and the third column shows the bias, which is the difference between the prediction result and the ground-truth. Theses visualizations validate our model’s ability to predict future weather states close to the ground-truth.
Figure 4: Visualization examples of future state prediction for Z500 compared with ground-truth.
Figure 5: Visualization examples of future state prediction for U10 compared with ground-truth.
5.4 Ablation studies
We analyze the importance of individual elements in EWMoE by removing one component at a time and observing the performance difference.
Effect of meteorology-specific embedding. We conduct experiments to compare the performance of EWMoE with and without meteorology-specific embedding to evaluate the its effectiveness in extracting the internal relations between meteorology variables. Figure 6(a) shows the superior performance of meteorology-specific embedding compared with the standard ViT embedding at all forecast time, indicating that it is a crucial component in capturing interactions between atmospheric variables through cross-attention.
(a) Effect of meteorology embedding.
(b) Effect of MoE layer.
Figure 6: Ablation studies showing the importance of each component in EWMoE. Similar trends are observed across different variables.
Effect of MoE layer. We evaluate the effectiveness of the MoE layer in EWMoE. As shown in Figure 6(b), EWMoE with the MoE layer significantly outperforms model with a standard feed-forward network, and the performance gap becomes larger as the forecast time increases. We attribute this result to the ability of the MoE layer, which allows increasing the model capacity and flexibility. We also note that EWMoE achieves this improvement without increasing computing requirements, as only a small portion of parameters are activated during training. This suggests that applying MoE structure to weather forecasting is promising. Moreover, we also conduct extensive experiments to evaluate the importance of auxiliary loss used in the MoE layer routing and the number of selective top-$k$ experts, as shown in Figure 7.
Figure 7: EWMoE improves consistently with auxiliary loss (left) and smaller $k$ (right).
6 Conclusion
In this paper, we introduce EWMoE, an advanced and effective deep learning model for weather forecasting. By integrating three novel components, meteorology-specific embedding, an MoE layer and two specific loss functions, it excels at a resolution of 0.25 and forecast time of up to 8 days, outperforming the leading models such as FourCastNet and ClimaX, and competing well with Pangu-Weather in short-range forecasting. It is worth mentioning that EWMoE achieves this superior performance with significantly less training data and computing resources, addressing the challenges of computational efficiency and prediction accuracy. Our study also provides insights for modeling the interactions among atmospheric variables, demonstrating the feasibility and potential of implementing the MoE paradigm in weather forecasting tasks. We hope that our work will inspire future work on applying effective MoE architecture to a wider range of climate researches.
[…]
Acknowledgements
The authors gratefully acknowledge the available of the ERA5 dataset on both pressure levels and single level provided by the European Centre for Medium-Range Weather Forecasts (ECMWF). Without their efforts in collecting, archiving, and disseminating the data, this work would not be feasible.
Declarations
  • Funding This work was supported by the National Natural Science Foundation of China (No. 62276047).
  • Conflict of interest The authors have no financial or non-financial interests to disclose.
  • Ethics approval and consent to participate The authors declare that this research did not require Ethics approval or Consent to participate since no experiments involving humans or animals have been conducted.
  • Consent for publication The authors of this manuscript all consent to its publication.
  • Data and code availability The code and data are available at https://github.com/Tomoyi/EWMoE.